OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 2y = –6 3x – 2y = 11 What is the solution to the system? a. (–1, –7) b. (1, –4) c. (2, –1) d. (3, –2)

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@Agent_Sniffles

OpenStudy (campbell_st):

go and download geoegebra its a free graphing package. http://www.geogebra.org/cms/ or the alternate method is to get some graph paper find the intercepts for each line, then graph the line to find the x- intercepts (where the line cute the x axis) let y = zero and solve for x for your 1st line if y = 0 2x +2*0 = -6 or 2x = -6 then x = -3 for y-interscepts let x = 0 and solve for y 2*0 + 2y = -6 2y = -6 you can finish... If you know where a line cuts the x and y axes you can draw the line. repeat the process for the other line... and your point if intersection will be obvious.

OpenStudy (mathstudent55):

Graph each equation and see where they intersect.

OpenStudy (anonymous):

solve for y then paste what you got for y in google seperated by commas 2x + 2y = –6, 3x – 2y = 11

OpenStudy (anonymous):

is it b? @timo86m

OpenStudy (anonymous):

x = 1, y = -4

OpenStudy (anonymous):

so yes you are right :D

OpenStudy (anonymous):

Yes, you got it right.